Mary and Joe Belinksy went out to dinner for Valentine‘s Day and paid for the meal using a credit card that has an APR of 20 percent. If they make only minimum payments each month, and do not repay the $100 cost of the dinner until 12 months later, how much did the dinner actually cost them?
Answer #1

Minimum Payments paid over 12 months = Interest Portion = Total Cost*APR = 100*20% = $20

Now, Assuming that full cost i.e. $100 is paid after 12 months,

Actual Cost of Dinner = Minimum Payments made+Actual Cost = 20+100 = $120